OpenStudy (anonymous):

1. Write the slope-intercept form of the line which passes through (4, 2) with a slope of 1/2. A. y = 1/2 x + 2 B. y = 1/2 x C. y = 1/2 x + 4

OpenStudy (anonymous):

y=1/2 x+c 2=2+c c=0 and so y=1/2x